[Hallc_running] [New Logentry] RC Daily Update

rmontgom at jlab.org rmontgom at jlab.org
Sun Nov 25 10:40:02 EST 2018

Logentry Text:
- Swing yesterday completed production data taking on Group 7 (Q2 = 3.98, x = 0.5, pi+).
- There was a ~1hr beam down time due to RF zone 1L24 dropping out of HV (DTM ID 5521) but other than that beam has overall been better (https://logbooks.jlab.org/entry/3631291). SHMS/HMS angles and momenta were changed for moving to Group 17 (Q2 = 5.5, x = 0.5). Data taking on this group also included an SHMS polarity change. Owl crew completed Group 17 and moved to 18.
- During owl shift the SHMS vacuum started dropping rapidly down to 1.55e-3 (https://logbooks.jlab.org/entry/3631178). This required an escorted access during which SHMS forepump was turned back on and SHMS turbo controller was reset (https://logbooks.jlab.org/entry/3631199). This was caused by the brownouts on 11/24. It would be good to discuss adding SHMS/HMS vacuum numbers into alarm handler, shift checklist etc and other checklist items before locking up swing shift Tuesday in case anything else has been affected by the brownouts.
- The day crew are currently completing Group 18 (Q2=5.5, x=0.5, z=0.6, pi+ and pi-), which will be followed by 3 (Q2=4, x=0.4, pi+) and 4 (Q2=4, x=0.4, pi-). The lowest z settings of groups 3 and 4 may be dropped due to time constraints.

RC Meeting:
- We have no planned RC meetings until Monday. Unless there is a special request.

Notices in effect for shift crews:
* NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''
* Shift leaders, please fill out the check list once per shift!
* It is no longer required to fill out the standard.kinematics file online, it is now done automatically
* Split the large desired charge between multiple runs (~1 hour per run)
* Listed currents are estimates. For runs at less than 70uA, increase current until SHMS singles rates are between 450k-500k.
* Remember to check trigger pre-scales before runs (PS6=0 all the time, adjusting PS4 to give at least 100Hz, and no more than 200Hz, of HMS singles, PS1=PS2=PS3=PS5=-1).
* Check that the pre-scaler gui is active and hasn't frozen/doesn't need re-started, especially between target changes
* How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Monday, Nov 26: Energy change to 1.86 GeV/pass and move to 500 MHz laser. Hall will go to restricted access Monday morning and aim is to be locked up again Tuesday Swing. There has been a separate e-mail/log entry about shift updates. So far these changes still stand but we will know more Tue/Wed about expected beam delivery - stay tuned for updates.

Planned accesses to the Hall:
Monday, Nov 26: Work in Hall should be coordinated with Walter K. Aerogel change by V. Berdnikov (~8 hours duration Mon). HGC work by B. Sawatzky and University of Regina team. D. Meekins will investigate target loop 3 JT valve issue. V. Kumar (Regina group) walkthrough Monday morning.
Tuesday, Nov 27: installation of brushes near the spectrometer wheels for S. Lassiter (originally planned for Tue/Wed).
Please let me know any other plans for work in the Hall Mon/Tue.

Special requests:
Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463). Has to be done before shut down in Dec.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3631299
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181125/f1c8e89f/attachment-0001.html>

More information about the Hallc_running mailing list